# Timer-XL: Long-Context Transformers for Unified Time Series Forecasting

### v2

@@ Results
As shown in Figure 3, the accuracy of univariate prediction can be generally improved by extending the daily context to monthly, thus emphasizing widespread context bottleneck in this field. We compare Timer-XL with PatchTST, where Timer-XL can surpass the state-of-the-art encoder-only Transformer when the context length persistently increases. It demonstrates that Timer-XL is proficient in handling extremely long contexts, and thus achieves breakthrough performance. It is also notable that the performance of monthly and yearly contexts improves slowly or even deteriorates, termed context oversatuation. It may stem from increased noise and training difficulty, which leaves a future direction to improve the context efficiency. Table 1 provide results on ERA5-S. Timer-XL consistently outperforms PatchTST on all sites, which can be credited to the maintenance of causality and token-wise supervision in the decoder-only architecture.
@@
Furthermore, we analyze the widespread non-stationary challenge in univariate tasks. It is commonly tackled by normalization (Kim et al., 2021) that greatly improves Transformer performance in previous benchmarks. However, we find it may be caused by the insufficient time span and training samples in these datasets. Thus, normalization enriches training samples by aligning time series with different means and variances to the same distribution. Instead, it makes Transformer constrained on the temporal variation within windows, preventing them from learning variations among windows and resulting in oversmooth predictions and failures in long contexts. In Table 1, we evaluate the performance on ERA5-S that includes sufficient samples. It validate the claim that long-context Transformers can achieve better results even without normalization.
@@
As shown in Tables 2-4 and Figure 4, Timer-XL achieves the best results on both previous and new benchmarks. Essentially, Transformers that explicitly capture inter-series dependencies, such as Moirai (Woo et al., 2024) and iTransformer, reasonably achieve decent performance in Table 2. Beyond iTransformer, Timer-XL can model fine-grained patch-wise temporal dependencies. With TimeAttention, Timer-XL outperforms Timer especially on high-dimensional time series ($13.2$% in ECL and $6.3$% in Traffic, with thousands of tokens in the context). Compared with the encoder-only Moirai, decoder-only Transformers excel at generalizing across varying prediction lengths in Table 3.
@@
Table 2: Multivariate forecasting ($96$-pred-$96$) of well-acknowledged benchmarks. Timer and Moirai are reproduced for evaluation. Results of other baselines are officially reported by Liu et al. (2023).
@@ 4.4 Pre-Trained Time Series Transformers
Pre-training aims to teach time-series Transformers transferable temporal dynamics. The outcome large time series model can better cope with widespread challenges of few-shot and zero-shot forecasting. To scale Timer-XL as large models, we conduct pre-training on UTSD (1 billion time points) from Liu et al. (2024c) for a domain-universal model, and our curated dataset ERA5-Large (40 years and 4920 stations) for a domain-specific model. On ERA5-Large, we evaluate in-dataset generalization performance of PatchTST and Timer-XL: pre-training on $80$% stations and $80$% time span and directly forecast on the remaining stations (variable generalization), time span (temporal generalization), and the cross-time and cross-station split. To evaluate out-of-dataset generalization performance, we perform zero-shot forecasting following Timer (2024c). Compared with Timer, the context length of Timer-XL is increased from $672$ to $2880$.
@@ Results
We provide in-dataset generalization performance of ERA5-Large in the middle of Figure 5. Timer-XL achieves better results than PatchTST in all cases, supporting that decoder-only architecture has stronger generalization performance. Figure 5 compares zero-shot forecasting results, where Timer-XL outperforms Timer on all benchmark datasets, indicating a promising direction to improve zero-shot forecasting by pre-training on a larger context.
@@ 4.5 Model Analysis
Ablation Study
As shown in Table 6, we conduct evaluations on TimeAttention to validate the effectiveness of position embeddings. As for variable embedding, the distinction between endogenous and exogenous variables can improve performance. Based on our observation of the learned $u>v$, we find that the token reasonably pays more attention to tokens of the endogenous variable. It leaves a prior to mask out minor dependencies that focuses less on exogenous variables. For the temporal dimension, other position embeddings are inferior to RoPE, since it uses the affine transformation, while others are additive, and thereby less confused with the same additive embedding for variables. As a counterpart Transformer, Moirai adopts the same combination of position embeddings, but its performance is still inferior to Timer-XL in Table 3 and 4, indicating that decoder-only Transformers are more compatible with such position embeddings.
Table 6: Embedding ablation in TimeAttention. For the temporal dimension, we compare prevalent relative and absolute position embeddings. For the variable dimension, we explore the effectiveness of the variable embedding that distinguishes endogenous and exogenous variables.
Design Temporal Variable Traffic Weather Solar-Energy ERA5-MS
MSE MAE MSE MAE MSE MAE MSE MAE
Ours RoPE (2024) with 0.340 0.238 0.157 0.205 0.162 0.221 0.164 0.307
Replace ALiBi (2021) with 0.351 0.246 0.162 0.212 0.188 0.210 0.167 0.308
Relative (2020) with 0.361 0.250 0.163 0.214 0.197 0.215 0.168 0.309
Absolute (2017) with 0.381 0.270 0.159 0.207 0.171 0.204 0.165 0.306
w/o RoPE (2024) w/o 0.361 0.254 0.171 0.217 0.181 0.221 0.235 0.373
w/o w/o 0.363 0.253 0.164 0.215 0.194 0.215 0.167 0.309
@@
Since the previous analysis of model efficiency on time-series Transformer predominantly focuses on the self-attention on 1D time series, we initially present a theoretical derivation of the computational complexity of Transformers on 2D time series, including the parameter counts, memory footprint, and FLOPs in Table 7. We find that other parts of Transformers, such as feed-forward network, have a complexity of $\mathcal{O}(NT)$ no matter which approach is adopted to handle multivariate time series. They also cause the dominant overhead in existing benchmarks, since their context length is not large enough, confirming our empirical results. Besides, we can introduce FlashAttention (Dao et al., 2022) to reduce the memory footprint and training speed, which is computationally equivalent and reduces the overall memory footprint of Timer-XL to $\mathcal{O}(NT)$ without affecting performance.
@@ A.2 FLOPs
For models that capture fine-grained intra- and inter-series dependencies (Timer-XL and Moirai) in multivariate series, $N$ is reflected as the enlarged number of tokens:
@@ B.2 Baseline Models
We aim to present Timer-XL as a foundation model for unified time series forecasting. We thoroughly include well-acknowledged and advanced models in each forecasting task. For univariate time series forecasting, we compare Timer-XL with PatchTST (Nie et al., 2022) under channel independence. For multivariate time series prediction, we report official results from Liu et al. (2023; 2024b); Ding et al. (2024), including UniRepLKNet (2024), iTransformer (2023), Corrformer (2023), DLinear (2023), TimesNet (2022), Non-stationary Transformer (2022b), Pyraformer (2021), Autoformer (2021) , StemGNN (2020), DeepAR (2020), and N-BEATS (2019). We further reproduce the performance of related Transformers: Timer (2024c) and Moirai (2024) based on their official repositories. For covariate-informed time series forecasting, we report the official results of TimeXer (2024b). For zero-shot forecasting, we follow Liu et al. (2024c) that predicts future length-$96$ windows in well-acknowledged datasets. In general, more than 20 baselines are included for a complete comparison.
@@ B.3 Implementation Details
All the experiments are implemented by PyTorch Paszke et al. (2019) on NVIDIA A100 Tensor Core GPUs. We employ the Adam optimizer (Kingma & Ba, 2014) and MSE loss for model optimization. We adopt channel independence from Nie et al. (2022) in univariate time series forecasting. Based on the prevalence of patch-level tokenization in the time series field, we reproduce typical Transformers: PatchTST (2022), Timer (2024c), and Moirai (2024) based on their official repositories, and keep their model hyperparameters and training configurations the same to evaluate the inherent capability of base models. The results of other baselines are based on the benchmark provided by Liu et al. (2023; 2024b); Ding et al. (2024); Wang et al. (2024b), which is fairly built on the configurations provided by their original paper. Detailed experimental configurations are provided in Table 10. We also report the standard deviations under three runs with different random seeds in Table 9, which exhibits that the performance of Timer-XL is stable.
@@
Timer 8 1024 128 8 96 0.00005 MSE 16384 10
@@ Appendix C Hyperparameter Sensitivity
Appendix D Supplementary Results
@@
Appendix E Limitations

### v3

@@ Results
As shown in Figure 3, the accuracy of univariate prediction can generally be improved by extending the daily context to monthly. We draw a similar conclusion on ERA5 (Table 14), where extending the context consistently helps in the specific model architecture. Empirically, Timer-XL, the generative Transformer, outperforms state-of-the-art encoder-only Transformer and linear forecaster in excessively long contexts. Further, we conduct representation analysis in Appendix E.4, revealing that Timer-XL is proficient at adaptively selecting information in vast observations, and thus achieves breakthrough performance. It is also noteworthy that the performance of monthly and yearly contexts improves slowly and deteriorates, which may stem from increased noise and training difficulty inherent in data, which leaves a future direction to improve the context efficiency. Table 1 provides results on ERA5-S. Timer-XL consistently outperforms PatchTST on all sites, which can be credited to the maintenance of causality and token-wise supervision in the decoder-only architecture.
@@
Furthermore, we analyze the widespread non-stationary challenge in univariate tasks. It is commonly tackled by normalization (Kim et al., 2021) that greatly improves Transformer performance in previous benchmarks. However, we find it may be caused by the insufficient time span and training samples in these datasets. Thus, normalization enriches training samples by aligning time series with different means and variances to the same distribution. Instead, it makes Transformer constrained on the temporal variation within windows, preventing them from learning variations among windows and resulting in oversmooth predictions and failures in long contexts. In Table 1 and Table 15, we evaluate the performance on ERA5 and widely-acknowledged datasets, which validates the claim that generative Transformers can achieve better results even without instance normalization.
@@
As shown in Tables 2-4 and Figure 4, Timer-XL achieves the best results on both previous and new benchmarks. Essentially, Transformers that explicitly capture inter-series dependencies, such as UniTST (Liu et al., 2024a) and iTransformer, reasonably achieve decent performance in Table 2. Beyond iTransformer, Timer-XL can model fine-grained patch-wise temporal dependencies. With TimeAttention, Timer-XL outperforms Timer especially on high-dimensional time series ($13.2$% in ECL and $6.3$% in Traffic, with thousands of tokens in the context). Compared with the encoder-only UniTST, decoder-only Transformers excel at generalizing across varying prediction lengths in Table 3.
@@
Table 2: Multivariate forecasting ($96$-pred-$96$) of well-acknowledged benchmarks. All models are trained from scratch. Results of baseline models are officially reported by Liu et al. (2023).
@@ 4.4 Pre-Trained Time Series Transformers
Pre-training enriches time-series Transformers with generalizable forecasting capabilities. The outcome large time series model can better cope with widespread challenges of few-shot and zero-shot forecasting. To scale Timer-XL as large models, we conduct pre-training on UTSD (Liu et al., 2024c) and LOTSA (Woo et al., 2024) for zero-shot forecasters. We also curate a large-scale multivariate dataset (ERA5-Large: 40 years and 4920 stations) for a domain-specific model. In this task, we evaluate in-dataset generalization performance of PatchTST and Timer-XL: pre-training on $80$% stations and $80$% time span and then forecast on the remaining stations (variable generalization), time span (temporal generalization), and the cross-time and cross-station split. To evaluate the benefit of pre-training with longer context, we compare the zero-shot performance of Timer (2024c) and Timer-XL by pre-training on UTSD (1B time points), where the context length of Timer-XL is increased from $672$ to $2880$. To establish a fair zero-shot benchmark, we pre-train Timer-XL on LOTSA (27B observations) such that baseline models are built upon a comparable pre-training scale.
@@ Results
We provide in-dataset generalization results of ERA5-Large in the middle of Figure 5 (a). Timer-XL achieves better results than PatchTST in all cases, supporting that decoder-only architecture has stronger generalization performance. Figure 5 (b) compares zero-shot performance of two generative Transformers pre-trained on the same UTSD, where Timer-XL outperforms previous Timer on all benchmark datasets, validating that long-context pre-training empowers large time series models. In Table 6, we provide a comprehensive zero-shot evaluation under a comparable pre-training scale and model size, where Timer-XL achieves notable performance with better sample efficiency. The versatility and scalability make it a promising backbone of foundation models.
@@
[…]
Models
[…]
$\textbf{Moirai}_{\textit{Small}}$
$\textbf{Moirai}_{\textit{Base}}$
$\textbf{Moirai}_{\textit{Large}}$
TimesFM
MOMENT
$\textbf{Chronos}_{\textit{Base}}$
$\textbf{Chronos}_{\textit{Large}}$
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTh1
0.394
0.413
0.363
0.390
0.401
0.402
0.376
0.392
0.381
0.388
0.414
0.404
0.688
0.557
0.440
0.393
0.441
0.390
ETTh2
0.280
0.343
0.287
0.344
0.297
0.336
0.294
0.330
0.296
0.330
0.315
0.349
0.342
0.396
0.308
0.343
0.320
0.345
ETTm1
0.621
0.512
0.308
0.355
0.418
0.392
0.363
0.356
0.380
0.361
0.361
0.370
0.654
0.527
0.454
0.408
0.457
0.403
ETTm2
0.204
0.292
0.192
0.281
0.214
0.288
0.205
0.273
0.211
0.274
0.202
0.270
0.260
0.335
0.199
0.274
0.197 0.271
ECL
0.141 0.236
0.135
0.229
0.212
0.300
0.162
0.249
0.155
0.237
0.291
0.355
0.154
0.231
0.152
0.229
Weather
0.167
0.224
0.169
0.223
0.198
0.222
0.220
0.217
0.199
0.211
0.243
0.255
0.203
0.238
0.194
0.235
$1^{\text{st}}$ Count
2
6
0
1
3
1
0
0
0
  • $\ast$ Traffic is not evaluated here because it is generally included in the pre-training datasets of large models.
@@ 4.5 Model Analysis
Since the previous analysis of model efficiency on time-series Transformer predominantly focuses on the self-attention on 1D time series, we initially present a theoretical derivation of the computational complexity of Transformers on 2D time series, including the parameter counts, memory footprint, and FLOPs in Table 7. We find that other parts of Transformers, such as feed-forward network, have a complexity of $\mathcal{O}(NT)$ no matter which approach is adopted to handle multivariate time series. They also cause the dominant overhead in existing benchmarks, since their context length is not large enough, confirming our empirical results. Further, we introduce FlashAttention (Dao et al., 2022) to reduce the memory footprint and training speed, which is computationally equivalent and reduces the overall memory footprint of Timer-XL to $\mathcal{O}(NT)$ without affecting performance.
@@ A.2 FLOPs
For models that capture fine-grained intra- and inter-series dependencies (Timer-XL and UniTST) in multivariate series, $N$ is reflected as the enlarged number of tokens:
@@ B.2 Baseline Models
We aim to present Timer-XL as a foundation model for unified time series forecasting. We thoroughly include well-acknowledged and advanced models in each forecasting task. For univariate time series forecasting, we compare Timer-XL with PatchTST (Nie et al., 2022) under channel independence. For multivariate time series prediction, we report official results from Liu et al. (2023; 2024b); Ding et al. (2024), including UniRepLKNet (2024), iTransformer (2023), Corrformer (2023), DLinear (2023), TimesNet (2022), Non-stationary Transformer (2022b), Pyraformer (2021), Autoformer (2021) , StemGNN (2020), DeepAR (2020), and N-BEATS (2019). We further reproduce the performance of related Transformers: Timer (2024c) and UniTST (2024a) based on their official repositories. For covariate-informed time series forecasting, we report the official results of TimeXer (2024b). For zero-shot forecasting, we follow Liu et al. (2024c) that predicts future length-$96$ windows in well-acknowledged datasets. Totally, more than 20 baselines are included for a complete comparison.
@@ B.3 Implementation Details
All the experiments are implemented by PyTorch (Paszke et al., 2019) on NVIDIA A100 Tensor Core GPUs. We employ the Adam optimizer (Kingma & Ba, 2014) and MSE loss for model optimization. We adopt channel independence from Nie et al. (2022) in univariate time series forecasting. Based on the prevalence of patch-level tokenization in the time series field, we reproduce typical Transformers: PatchTST (2022), Timer (2024c), and UniTST (2024a) based on their official repositories, and keep their model hyperparameters and training configurations the same to evaluate the inherent capability of base models. The results of other baselines are based on the benchmark provided by Liu et al. (2023; 2024b); Ding et al. (2024); Wang et al. (2024b), which is fairly built on the configurations provided by their original paper. Detailed experimental configurations are provided in Table 10. We also report the standard deviations under three runs with different random seeds in Table 9, which exhibits that the performance of Timer-XL is stable.
@@
Timer (Liu et al., 2024c) 8 1024 128 8 96 0.00005 MSE 16384 10
Timer-XL 8 1024 128 8 96 0.001 MSE 32768 -
$\text{Moirai}_{\text{Small}}$ LOTSA 6 384 64 6 -
$\text{Moirai}_{\text{Base}}$ (Woo et al., 2024) 12 768 64 12 - Woo et al. (2024)
$\text{Moirai}_{\text{Large}}$ 24 1024 64 16 -
@@ Appendix C Hyperparameter Sensitivity
Appendix D Showcases
To facilitate a clear comparison among various models, we present additional prediction visualization from diverse datasets in Figure 9 and 10. Showcases are randomly selected from Timer-XL and the following time-series Transformers: PatchTST (2022), Timer (2024c), and UniTST (2024a). Among them, Timer-XL presents the most accurate predictions.
Figure 9: Visualization results on univariate time series dataset. We adopt the forecasting setting of $2880$-pred-$96$ on ECL, ETTh1 and Traffic, and $2016$-pred-$96$ on PEMS.
Figure 10: Visualization results on multivariate time series dataset. We adopt the forecasting setting of $672$-pred-$96$ on ETTh1 (7 Variables) and Traffic (862 Variables).
Appendix E Supplementary Results
E.1 Full Result of Multivariate Forecasting
@@
E.2 Full Result of Zero-Shot Forecasting
Table 12 provides the full results of zero-shot forecasting on the benchmark from Wu et al. (2022). We build Timer-XL based on the configuration in Table 10, which is pre-trained on the aggregated datasets of UTSD (Liu et al., 2024c) and LOTSA (Woo et al., 2024). The patch size of Timer-XL is set as $96$ and we conduct rolling forecast to obtain the desired forecast length in $\{96,192,336,720\}$.
We evaluate most advanced large models based on their official model checkpoints, including Time-MoE (Shi et al., 2024), Moirai (Woo et al., 2024), TimesFM (Das et al., 2023), MOMENT Goswami et al. (2024), and Chronos (Ansari et al., 2024). We conduct zero-shot evaluations on datasets that are not included during the pre-training of corresponding models. For each of the evaluated model, we use their maximum input length during inference. The metric (MSE/MAE) is averaged from all predicted windows in the test split.
Table 12: Full results of zero-shot forecasting (pred-$\{96,192,336,720\}$). We add new baseline large models, such as Time-MoE (Shi et al., 2024). Our model size is comparable with $\textbf{Moirai}_{\textit{Base}}$, which is pre-trained on UTSD (Liu et al., 2024c) and LOTSA (Woo et al., 2024).
Models
[…]
$\textbf{Moirai}_{\textit{Small}}$
$\textbf{Moirai}_{\textit{Base}}$
$\textbf{Moirai}_{\textit{Large}}$
TimesFM
MOMENT
$\textbf{Chronos}_{\textit{Base}}$
$\textbf{Chronos}_{\textit{Large}}$
Metric
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.317
0.356
0.338
0.368
0.309
0.357
0.281
0.341
0.418
0.392
0.363
0.356
0.380
0.361
0.361
0.370
0.654
0.527
0.454
0.408
0.457
0.403
192
0.358
0.381
0.353
0.388
0.346
0.381
0.305
0.358
0.431
0.405
0.388
0.375
0.412
0.383
0.414
0.405
0.662
0.532
0.567
0.477
0.530
0.450
336
0.386
0.401
0.381
0.413
0.373
0.408
0.369
0.395
0.433
0.412
0.416
0.392
0.436
0.400
0.445
0.429
0.672
0.537
0.662
0.525
0.577
0.481
720
0.430
0.431
0.504
0.493
0.475
0.477
0.469
0.472
0.462
0.432
0.460
0.418
0.462
0.420
0.512
0.471
0.692
0.551
0.900
0.591
0.660
0.526
Avg
0.373
0.392
0.394
0.415
0.376
0.405
0.356 0.391
0.436
0.410
0.406
0.385
0.422
0.391
0.433
0.418
0.670
0.536
0.645
0.500
0.555
0.465
ETTm2
96
0.189
0.277
0.201
0.291
0.197
0.286
0.198
0.288
0.214
0.288
0.205
0.273
0.211
0.274
0.202
0.270
0.260
0.335
0.199
0.274
0.197 0.271
192
0.241
0.315
0.258
0.334
0.250
0.322
0.235
0.312
0.284
0.332
0.275
0.316
0.281
0.318
0.289
0.321
0.289
0.350
0.261
0.322
0.254
0.314
336
0.286
0.348
0.324
0.373
0.337
0.375
0.293
0.348
0.331
0.362
0.329
0.350
0.341
0.355
0.360
0.366
0.324
0.369
0.326
0.366
0.313
0.353
720
0.375
0.402
0.488
0.464
0.480
0.461
0.427
0.428
0.402
0.408
0.437
0.411
0.485
0.428
0.462
0.430
0.394
0.409
0.455
0.439
0.416
0.415
Avg
0.273
0.336
0.317
0.365
0.316
0.361
0.288
0.344
0.307
0.347
0.311
0.337
0.329
0.343
0.328
0.346
0.316
0.365
0.310
0.350
0.295
0.338
ETTh1
96
0.369
0.391
0.357
0.381 0.350
0.382
0.349
0.379
0.401
0.402
0.376
0.392
0.381
0.388
0.414
0.404
0.688
0.557
0.440
0.393
0.441
0.390
192
0.405
0.413
0.384
0.404
0.388 0.412
0.395
0.413
0.435
0.421
0.412
0.413
0.434
0.415
0.465
0.434
0.688
0.560
0.492
0.426
0.502
0.524
336
0.418
0.423
0.411
0.434
0.411
0.430
0.447
0.453
0.438
0.434
0.433
0.428
0.485
0.445
0.503
0.456
0.675
0.563
0.550
0.462
0.576
0.467
720
0.423
0.441
0.449
0.477
0.427
0.455
0.457
0.462
0.439
0.454
0.447
0.444
0.611
0.510
0.511
0.481
0.683
0.585
0.882
0.591
0.835
0.583
Avg
0.404
0.417
0.400
0.424
0.394
0.419
0.412
0.426
0.428
0.427
0.417
0.419
0.480
0.439
0.473
0.443
0.683
0.566
0.591
0.468
0.588
0.466
ETTh2
96
0.283
0.342
0.305
0.359
0.302
0.354
0.292
0.352
0.297
0.336
0.294
0.330
0.296
0.330
0.315
0.349
0.342
0.396
0.308
0.343
0.320
0.345
192
0.340
0.379
0.351
0.386
0.364
0.385
0.347
0.379
0.368
0.381
0.365
0.375
0.361
0.371
0.388
0.395
0.354
0.402
0.384
0.392
0.406
0.399
336
0.366
0.400
0.391
0.418
0.417
0.425
0.406
0.419
0.370
0.393
0.376
0.390
0.390
0.390
0.422
0.427
0.356
0.407
0.429
0.430
0.492
0.453
720
0.397
0.431
0.419
0.454
0.537
0.496
0.439
0.447
0.411
0.426
0.416
0.433
0.423
0.418
0.443
0.454
0.395
0.434
0.501
0.477
0.603
0.511
Avg
0.347
0.388
0.366
0.404
0.405
0.415
0.371
0.399
0.361
0.384
0.362
0.382
0.367
0.377
0.392
0.406
0.361
0.409
0.405
0.410
0.455
0.427
ECL
96
0.141
0.237
0.189
0.280
0.160
0.250
0.153
0.241
0.745
0.680
0.154
0.231 0.152
0.229
192
0.159
0.254
0.205
0.292
0.175
0.263
0.169
0.255
0.755
0.683
0.179
0.254
0.172
0.250
336
0.177
0.272
0.221
0.307
0.187
0.277
0.187
0.273
0.766
0.687
0.214
0.284
0.203
0.276
720
0.219
0.308
0.258
0.335
0.228 0.309
0.237
0.313
0.794
0.696
0.311
0.346
0.289
0.337
Avg
0.174
0.278
0.218
0.303
0.187
0.274
0.186
0.270
0.765
0.686
0.214
0.278
0.204
0.273
Weather
96
0.171
0.225
0.160
0.214
0.159 0.213
0.157
0.211
0.198
0.222
0.220
0.217
0.199
0.211
0.243
0.255
0.203
0.238
0.194
0.235
192
0.221
0.271
0.210
0.260
0.215
0.266
0.208
0.256
0.247
0.265
0.271
0.259
0.246
0.251
0.278
0.329
0.256
0.290
0.249
0.285
336
0.274
0.311
0.274
0.309
0.291
0.322
0.255
0.290
0.283
0.303
0.286
0.297
0.274 0.291
0.306
0.346
0.314
0.336
0.302
0.327
720
0.356
0.370
0.418
0.405
0.415
0.400
0.405
0.397
0.373
0.354
0.373
0.354
0.337
0.340
0.350
0.374
0.397
0.396
0.372
0.378
Avg
0.256
0.294
0.265
0.297
0.270
0.300
0.256
0.288
0.275
0.286
0.287
0.281
0.264
0.273
0.294
0.326
0.292
0.315
0.279
0.306
$1^{\text{st}}$ Count
15
10
2
1
3
0
10
7
0
0
0
5
1
10
0
1
2
0
0
0
0
2
  • $\ast$ Dataset for pre-training is not evaluated on corresponding models, which is denoted by a dash (--).
  • $\ast$ Traffic from Wu et al. (2022) is generally used during the pre-training of large models and thus not evaluated here.
  • $\ast$ Our model checkpoint is available at https://huggingface.co/thuml/timer-base-84m.
E.3 Ablation Study of TimeAttention
We conduct evaluations on TimeAttention to validate the effectiveness of position embeddings. As for variable embedding, the distinction between endogenous and exogenous variables can improve performance. Based on our observation of the learned $u>v$, we find that the token reasonably pays more attention to tokens of the endogenous variable. It leaves a prior to mask out minor dependencies that focuses less on exogenous variables. For the temporal dimension, other position embeddings are inferior to RoPE, since it uses the affine transformation, while others are additive, and thereby less confused with the same additive embedding for variables.
Table 13: Embedding ablation in TimeAttention. For the temporal dimension, we compare prevalent relative and absolute position embeddings. As for the variable dimension, we explore the effectiveness of the variable embedding that distinguishes endogenous and exogenous variables.
Design Temporal Variable Traffic Weather Solar-Energy ERA5-MS
MSE MAE MSE MAE MSE MAE MSE MAE
Timer-XL RoPE (2024) with 0.340 0.238 0.157 0.205 0.162 0.221 0.164 0.307
Replace ALiBi (2021) with 0.351 0.246 0.162 0.212 0.188 0.210 0.167 0.308
Relative (2020) with 0.361 0.250 0.163 0.214 0.197 0.215 0.168 0.309
Absolute (2017) with 0.381 0.270 0.159 0.207 0.171 0.204 0.165 0.306
w/o RoPE (2024) w/o 0.361 0.254 0.171 0.217 0.181 0.221 0.235 0.373
w/o w/o 0.363 0.253 0.164 0.215 0.194 0.215 0.167 0.309
E.4 Supplementary Results of Long-Context Forecasting
Long context is a basic indicator of foundation models, which can support emergence capabilities such as prompting, in-context learning, retrieval-augmented generation, etc. However, the long-context forecasting paradigm receives less attention in the current community, which can be due to the lack of benchmarks. In the meteorological ERA5, it is necessary to support the context of more than years to contain a specific cycle (such as El Nino). In Table 14, the performance of Timer-XL and DLinear generally improves with the increased context length. By contrast, it reveals the performance degradation of PatchTST. Similar to the observations in Figure 3, the encoder-only architecture produces inferior predictions after thousands of time points, which can be concealed due to the short context adopted in previous benchmarks. Although PatchTST has conducted an initial exploration in the context of hundreds of time points, it inappropriately works in ever-long contexts. Therefore, we believe that context bottlenecks deserve further exploration in this community.
Table 14: Performance on ERA5 (pred-1day). Lookback lengths vary from daily to yearly contexts.
      Models       Timer-XL       PatchTST       DLinear
      Metric       MSE |||| MAE       MSE |||| MAE       MSE |||| MAE
      Lookback-8 (1 Day)       0.0847 |||| 0.2100       0.0897 |||| 0.2196       0.0970 |||| 0.2276
      Lookback-32 (4 Day)       0.0713 |||| 0.1928       0.0778 |||| 0.2080       0.0841 |||| 0.2113
      Lookback-56 (1 Week)       0.0688 |||| 0.1891       0.0785 |||| 0.2082       0.0814 |||| 0.2081
      Lookback-224 (1 Month)       0.0675 |||| 0.1868       0.0745 |||| 0.2042       0.0788 |||| 0.2048
      Lookback-960 (4 Month)       0.0667 |||| 0.1863       0.1194 |||| 0.2696       0.0773 |||| 0.2031
      Lookback-2944 (1 Year)       0.0663 |||| 0.1857       0.1109 |||| 0.2638       0.0763 |||| 0.2024
Representation Analysis
We further delve into long-context modeling from the perspective of learned representations. As shown in Figure 11, the decoder-only model can selectively focus on the previous context while PatchTST wrongly focuses on noisy parts. Since causality is the basis of forecasting, using causal masks leads to coherent token embeddings, while the unmasked attention mechanism may break the causality and prevent the model from telling each tokens.
Figure 11: Case studies of learned attention in encoder-/decoder-only Transformers.
Normalization
Section 4.1 has discussed instance normalization (Kim et al., 2021). It generally improves the performance of the previous encoder-only Transformers but leads to special problems in generative Transformers (e.g., unmatched statistics in multi-step autoregression). However, it is indicative that Timer-XL without ReVIN can achieve competitive performance on well-acknowledged benchmarks in Table 15, while the performance of PatchTST may heavily rely on this normalization.
Table 15: Evaluations ($672$-pred-$96$) on the effect of ReVIN (Kim et al., 2021) on Transformers.
Models Timer-XL with ReVIN Timer-XL w/o ReVIN PatchTST with ReVIN PatchTST w/o ReVIN
Metric MSE |||| MAE MSE |||| MAE MSE |||| MAE MSE |||| MAE
ETTh1 0.364 |||| 0.397 0.370 |||| 0.401 0.370 |||| 0.399 0.421 |||| 0.448
Weather 0.157 |||| 0.205 0.151 |||| 0.205 0.149 |||| 0.198 0.173 |||| 0.242
ECL 0.127 |||| 0.219 0.130 |||| 0.225 0.129 |||| 0.222 0.138 |||| 0.244
E.5 Illustration of TimeAttention
Although the formulation to generalize from 1D sequences to multivariate time series is straightforward, Timer-XL is built on a generative Transformer, an underexploited backbone among current time series models. As shown in Figure 12, challenges lie in capturing fine-grained dependencies between all variables in the patch level, while maintaining temporal causality in multiple sequences. Technically, we introduce the masking formulation, whose key lies in the grouped causality of flattened 2D sequences. We derive it based on the Kronecker Product, which disentangles the large attention map into formalizable temporal and variable dependencies. It can be naturally extended to covariates or pre-defined variable dependencies, which may inspire a lot of future explorations.
Figure 12: Illustration of TimeAttention for modeling univariate and multivariate time series.
Appendix F Limitations